# TRAWL: Tensor Reduced and Approximated Weights for Large Language Models

### v2

@@ Abstract
Recent research has shown that pruning large-scale language models for inference is an effective approach to improving model efficiency, significantly reducing model weights with minimal impact on performance. Interestingly, pruning can sometimes even enhance accuracy by removing noise that accumulates during training, particularly through matrix decompositions. However, recent work has primarily focused on single matrix decompositions or lower precision techniques, which may fail to fully capture structural patterns. To address these limitations, we introduce TRAWL (Tensor Reduced and Approximated Weights for Large Language Models), a technique that applies tensor decomposition across multiple weight matrices to effectively denoise LLMs by capturing global structural patterns. Our experiments show that TRAWL improves model performance by up to 16% over baseline models on benchmark datasets, without requiring additional data, training, or fine-tuning.
TRAWL: Tensor Reduced and Approximated Weights
for Large Language Models
Yiran Luo1*, Het Patel1*, Yu Fu1, Dawon Ahn1, Jia Chen2, Yue Dong1, Evangelos E. Papalexakis1 1 Department of Computer Science and Engineering, UC Riverside, USA 1 {yluo147, hpate061, yfu093, dahn017, yued, epapalex}@ucr.edu 2 Department of Electrical and Computer Engineering, UC Riverside, USA 2 jiac@ucr.edu * These authors contributed equally to this work.
1 Introduction
The success of Transformer-based models like the GPT series Radford et al. (2018, 2019); Brown et al. (2020), and LLaMA Touvron et al. (2023) is often attributed to their large scale Kaplan et al. (2020). While the large number of parameters allows these models to capture complex patterns, training and deploying them requires significant computational resources and energy Samsi et al. (2023); Zhao et al. (2022), making them impractical for many real-world applications.
To address these limitations, recent studies have explored post-training compression techniques, showing minimal impact on inference performance Han et al. (2015); Sharma et al. (2023); Saha et al. (2024). Interestingly, some research suggests that certain pruning methods can reduce computation time while also enhancing performance. For instance, LASER Sharma et al. (2023) applies Singular Value Decomposition (SVD) on weight matrices for model compression, demonstrating that removing noise introduced during training can improve performance. CALDERA Saha et al. (2024) builds on LASER by further compressing weight matrices into lower precision formats, achieving significant size reductions and lower computational costs. However, both methods operate on individual matrices, potentially missing broader structural patterns across the model.
@@ 1 Introduction
Our experiments on two Transformer-based models- RoBERTa Liu et al. (2019) and GPT-J Wang and Komatsuzaki (2021) across three datasets (BigBench WikiQA Srivastava et al. (2022), BiosProfession De-Arteaga et al. (2019), and HotpotQA Yang et al. (2018)) demonstrate that TRAWL consistently outperforms single matrix decomposition approaches, with accuracy improvements of up to 16%. By leveraging advanced tensor decomposition, TRAWL effectively reduces noise introduced during training and enhances performance without requiring additional training or fine-tuning. These results highlight TRAWL as a practical and efficient optimization technique for LLMs, delivering significant accuracy gains over baseline models.
@@
We conducted extensive experiments on two models across three datasets, showing that TRAWL improves accuracy and generalization over existing methods without additional training or fine-tuning, making it highly practical for real-world applications. All code from this study will be publicly available upon publication to support transparency and future research.111https://anonymous.4open.science/r/TRAWL
@@ 2 Related Work
The two most relevant works to ours are LAyer SElective Rank Reduction (LASER) Sharma et al. (2023) and CALDERA Saha et al. (2024), which leverage matrix decomposition and low-rank approximation to compress LLMs. LASER applies singular value decomposition (SVD) to reduce weight matrix rank, improving performance by removing noise. CALDERA combines low-precision and low-rank matrices to reduce memory usage while maintaining accuracy.
Other tensor-based methods, such as Low-Rank Economic Tensor-Train Adaptation (LoRETTA) Yang et al. (2024) significantly reduce trainable parameters through tensor-train decomposition, while techniques like Tensor Contraction Layers (TCLs) Kossaifi et al. (2017) incorporate tensor contractions as end-to-end trainable neural network layers to reduce the dimensionality of the activation tensors offering a way to reduce model parameters.
Building on these approaches, TRAWL fully leverages tensor structures in LLMs by applying tensor decomposition across multiple weight matrices, achieving performance improvements without additional training or fine-tuning.
@@ 3 Methodology
This section provides background on tensor decomposition and describes our proposed method.
@@ 3.1 Tensor Decomposition
Tensor decomposition is a technique for analyzing multi-dimensional data by breaking a tensor into lower-dimensional components Sidiropoulos et al. (2017). A tensor, essentially a multi-dimensional array, can be decomposed using various tensor decomposition methods. In this paper we focus on two: CANDECOMP/PARAFAC (CP) decomposition and Tucker decomposition.
@@ 3.2 Proposed Approach
TRAWL focuses on tensor compression across multiple weight matrices, distinguishing itself from existing methods Sharma et al. (2023); Saha et al. (2024) by emphasizing how matrices are stacked for decomposition.
We propose an effective approach by stacking and decomposing weight matrices one layer at a time. In more details, we stack the selected matrices for each layer individually and apply tensor decomposition layer by layer. Specifically, for each layer , we form a 3-mode tensor :
Tensor decomposition is applied to each separately, and the impact on performance is evaluated layer by layer. The matrices in each layer are replaced by their decomposed counterparts and reverted after evaluation. This process is repeated for all layers. The approach is depicted in Figure 1.
Additionally, we explored two conventional tensor formulation methods (Appendix 7.2): stacking selected matrices across all layers and stacking by dividing the matrices into three-layer segments. Our results demonstrate that both approaches underperform compared to our one-layer-at-a-time strategy. A detailed experimental comparison is provided in Appendix 7.3.
@@ 4.1 Experimental Setups
We primarily focus on two models, GPTJ-6B Wang and Komatsuzaki (2021) and RoBERTa Liu et al. (2019), and tested them with three datasets, BigBench WikiQA Srivastava et al. (2022), BiosProfession De-Arteaga et al. (2019), and HotPotQA Yang et al. (2018). In our experiments, we applied two tensor decomposition methods—CP decomposition and Tucker decomposition—to better understand the effect of tensor decomposition on LLMs. These were compared with baseline experiments where no tensor decomposition was applied.
To apply tensor decomposition, we transpose the Fully Connected (FC) input or output matrix to align their shapes for stacking and decomposition. After decomposition and low-rank approximation, the matrix is transposed back to its original shape before being substituted into the model. For the Query, Key, Value, and Output from attention heads (QKVO) matrices, stacking is straightforward since they share the same shape, allowing easy decomposition. More details are in Appendix 7.1.
@@ 4.2 Results
Table 1 compares the performance of the baseline, LASER, and TRAWL methods for both RoBERTa and GPT-J models across three datasets: BigBench WikiQA, BiosProfession, and HotpotQA. Additional results from other tensorization strategies, including segmented layers and across-the-model experiments, are provided in Appendix 7.3.
First, we observe that TRAWL, our tensor decomposition approach across multiple weight matrices, consistently outperforms the baseline compression method, LASER. For the BigBench WikiQA dataset, TRAWLCP delivers the best accuracy, with 43.46% for RoBERTa and 68.1% for GPT-J, outperforming LASER and the baseline. It also achieves the lowest loss for GPT-J, while LASER has the lowest for RoBERTa. In the BiosProfession dataset, TRAWLCP continues to lead with 73.07% accuracy for RoBERTa and 82.35% for GPT-J, while improving GPT-J’s loss. For HotpotQA, the improvements are smaller, with TRAWLCP slightly ahead of LASER for GPT-J, and similar results across methods for RoBERTa.
In summary, TRAWL consistently outperforms the baseline and LASER methods, particularly with CP decomposition in the FC layers. The most significant improvements are seen with the GPT-J model, especially on the BigBench WikiQA and BiosProfession datasets. The observed benefits, however, are dataset-dependent, with HotpotQA showing only marginal improvements.
@@ 4.3 Ablations
To evaluate the effect of stacking and decomposing different layers, we compared the performance of CP-decomposed fully connected (FC) weight matrices across various layers of RoBERTa on multiple datasets. The most notable improvements occurred in the FC weight matrices of the final layers, where tensor decomposition led to significant accuracy gains. In contrast, decomposing the QKVO matrices yielded minimal benefits and, in some cases, even reduced performance. As shown in Figure 2, the later-stage FC layers exhibited the largest accuracy improvements, while earlier layers, particularly the QKVO matrices, showed little to no benefit, and in some instances, decreased accuracy. We also observe similar trends on GPT-J, where the results are presented in Appendix Figure 4.
@@
Last but not least, the one-layer-at-a-time approach usually resulted in the highest performance gains near the final few layers when decomposing FC weights. However, no significant improvements were observed for QKVO matrices. These results align with the findings of Sharma et al. (2023), highlighting that FC layers are key targets for effective compression. In this work we demonstrate that there exists meaningful low-rank structure when treating the FC weights jointly as a tensor. In the future, we aim to explore whether more complex models or alternative methods can yield improvements for QKVO matrices.
@@ 5 Conclusion
In this study, we introduce TRAWL, a novel method for denoising LLMs using tensor decomposition. TRAWL extracts specific weight matrices, stacks them into a tensor, and applies decomposition to replace the original matrices with low-rank approximations. Our results show that layer-by-layer decomposition and selective layer stacking significantly improve performance, with the most notable gains seen in the fully connected layers of the terminal layers. These findings highlight the effectiveness of adaptive decomposition techniques in enhancing LLM efficiency, sometimes even outperforming the original models through denoising. Future work will explore alternative decomposition methods, refine rank selection, and expand TRAWL’s application to a broader range of models and tasks, driving more efficient and accessible AI systems.

### v3

@@ Abstract
Recent research has shown that factorizing large-scale language models for inference is an effective approach to improving model efficiency, significantly reducing model weights with minimal impact on performance. Interestingly, factorization can sometimes even improve accuracy by removing the noise that accumulates during training, particularly through matrix decompositions. However, recent work has primarily focused on single-matrix decompositions or lower precision techniques, which may fail to fully capture structural patterns. To address these limitations, we introduce TRAWL (Tensor Reduced and Approximated Weights for Large Language Models), a technique that applies tensor decomposition across multiple weight matrices to effectively denoise LLMs by capturing both global and local structural patterns. Our experiments show that TRAWL improves the model performance by up to 16% over baseline models on benchmark datasets, without requiring additional data, training, or fine-tuning.
Keywords:
LLM Compression Tensor Decomposition Low-rank Decomposition Denoising1 Introduction
The success of Transformer-based models like the GPT series [11, 12, 1], and LLaMA [19] is often attributed to their large scale [5]. While the large number of parameters allows these models to capture complex patterns, training and deploying them requires significant computational resources and energy [14, 24], making them impractical for many real-world applications.
To address these limitations, recent studies have explored post-training compression techniques, showing minimal impact on inference performance [3, 15, 13]. Interestingly, some research suggests that certain factorization methods can reduce computation time while also enhancing performance. For instance, LASER [15] applies Singular Value Decomposition (SVD) on weight matrices for model compression, demonstrating that removing noise introduced during training can improve performance. CALDERA [13] builds on LASER by further compressing weight matrices into lower precision formats, achieving significant size reductions and lower computational costs. However, both methods operate on individual matrices, potentially missing broader structural patterns across the model.
@@
Our experiments on two Transformer-based models—RoBERTa [9] and GPT-J [20] across three datasets (BigBench WikiQA [18], BiosProfession [2], and HotpotQA [22]) demonstrate that TRAWL consistently outperforms single matrix decomposition approaches, achieving accuracy improvements of up to 16%. By leveraging advanced tensor decomposition, TRAWL effectively reduces noise introduced during training and enhances performance without requiring additional training or fine-tuning. These results highlight TRAWL as a practical and efficient optimization technique for LLMs, delivering significant accuracy gains over baseline models.
@@
We conducted extensive experiments on two models across three datasets, showing that TRAWL improves accuracy and generalization over existing methods without additional training or fine-tuning, making it highly practical for real-world applications. All code from this study is publicly available to support transparency and further research at GitHub Link.
@@ 2 Related Work
Recent efforts in neural network compression have focused on matrix or tensor decomposition techniques to reduce computational or storage requirements while maintaining or even improving performance.
LAyer SElective Rank reduction (LASER) [15] and Calibration Aware Low-Precision DEcomposition with Low-Rank Adaptation (CALDERA) [13] are two representative methods which utilize matrix decomposition and low-rank approximation on LLM weights. LASER reduces the rank of weight matrices with singular value decomposition (SVD), enhancing performance by reducing noise from higher-order components. CALDERA decomposes a weight matrix into a low-precision backbone matrix and low-rank factors, reducing the model’s memory footprint while maintaining performance.
A recent study [23] challenges the assumption that the model weights in transformers are low rank. Their study, titled "Compressing Transformers: Features Are Low-Rank, but Weights Are Not!" demonstrates that while the features (i.e., activations) of transformers exhibit low-rank characteristics, the model weights do not. This finding is particularly significant, as it contrasts with the foundation of several compression techniques, such as LASER, which assume that low-rank weight approximations can be beneficial. The authors propose an adaptive approach to compress model features instead of weights, achieving substantial parameter reductions with minimal accuracy loss using a few-shot unsupervised method.
We acknowledge this distinction, but also observe that, in some cases, low-rank approximations of weights, although not necessarily as low-rank as one would normally expect from the popular use of the term, can still offer performance improvements with approximation ranks that are still lower than the original dimensionality of the weights. This nuanced understanding aligns with our results, where we find that certain configurations and ranks, even if not conventionally low (low compared to the dimensionality of weights), provide significant improvements. Thus, while our approach involves weight matrix decomposition, it does so with a flexible consideration of rank, corroborating, and extending the findings of the "Compressing Transformer" study.
Tensor train decomposition (TTD) [10] decomposes weight tensors into low-rank chains, effectively reducing parameters and computational complexity. Similarly, tensor ring decomposition (TRD) [25] arranges tensors in a ring structure for further model size reduction.
Lebedev et al. [8] proposed a two-step approach for speeding up the training for large convolutional neural networks (CNNs) using tensor decomposition and discriminative fine-tuning, which achieves significant speed-ups in training with minor accuracy drops.
Tensor Contraction Layers (TCLs) [6] incorporated tensor contractions as end-to-end trainable neural network layers, leading to significant model compression with minimal impact on accuracy and, in some cases, improved performance.
Low-Rank Adaptation (LoRA) [4] focuses on introducing trainable low-rank matrices into the transformer architecture for efficient fine-tuning, significantly reducing trainable parameters, and lowering GPU memory requirements without sacrificing model performance. Similarly, Low-Rank Economic Tensor-Train Adaptation (LoRETTA) [21] employs tensor-train decomposition to achieve significant parameter reduction while maintaining performance.
Our method, TRAWL, builds on these approaches by fully exploiting tensor structures in LLMs, achieving enhanced performance without additional training or fine-tuning. To our knowledge, this study is the first to explore the performance benefits of tensor factorization in LLMs in a setting where fine-tuning does not occur.
@@ 3 Methodology
This section provides background information on tensor decomposition and describes our proposed method.
@@ 3.1 Tensor Decomposition
Tensor decomposition is a technique for analyzing multidimensional data by breaking a tensor into lower dimensional components [17]. A tensor, essentially a multidimensional array, can be decomposed using various tensor decomposition methods. In this paper, we focus on two: CANDECOMP/PARAFAC (CP) decomposition and Tucker decomposition.
@@ 3.2 Proposed Approach
TRAWL focuses on tensor compression across multiple weight matrices, distinguishing itself from existing methods [15, 13] by emphasizing how matrices are stacked for decomposition.
In our approach, we explore different methods of tensor formulation to effectively apply tensor decomposition across multiple weight matrices within the transformer architecture. Specifically, we examine:
3.2.1 One Layer at a Time
In this setting, we stack the selected matrices for each layer individually and apply tensor decomposition layer by layer. Specifically, for each layer , we form a 3-mode tensor :
Tensor decomposition is applied to each layer, denoted as , separately. The impact on performance is evaluated layer by layer. In this process, the matrices of each layer are replaced with their decomposed and reconstructed versions. After evaluation, the original matrices are restored. This procedure is repeated for all layers. The approach is illustrated in Figure 1.
3.2.2 Across the Entire Model
In this setting, we stack the selected matrices (either QKVO or FC) from all layers of the model into a single higher-order tensor. This global stacking enables tensor decomposition across the entire set of matrices, capturing interactions across all layers. Formally, let represent the weight matrices of layer , where for the attention layers, or for the FC layers. We then construct a 3-mode tensor :
where is the number of layers. Tensor decomposition is applied to to capture global interactions across layers. This is illustrated in Figure 2.
3.2.3 Segmented Layers (Early, Middle, Last)
In this configuration, the layers are split into three segments: the initial third, intermediate third, and final third. Each segment undergoes separate tensor decomposition and performance evaluation, allowing analysis of the technique’s impact across sections of the model architecture to determine which segments benefit most.
@@ 4.1 Experimental Setups
We primarily focus on two models, GPTJ-6B [20] and RoBERTa [9], and tested them with three datasets, BigBench WikiQA [18], BiosProfession [2], and HotPotQA [22]. In our experiments, we applied two tensor decomposition methods—CP decomposition and Tucker decomposition—to better understand the effect of tensor decomposition on LLMs. These were compared with baseline experiments where no tensor decomposition was applied.
To effectively utilize tensor decomposition and low-rank approximation, we transpose one of the FC weight matrices when stacking them to form a tensor. This is necessary because the FC input and FC output matrices are transposed relative to each other in shape. By transposing, we can appropriately structure it as a suitable tensor for decomposition. Once the tensor decomposition is performed and a low-rank approximation is obtained, we transpose the resulting matrix back to its original shape. This transformed matrix then substitutes the original matrix in the LLM, ensuring that the model’s architecture remains consistent while benefiting from the reduced complexity and enhanced efficiency provided by the tensor decomposition process. For the QKVO matrices, we simply stack them together to form the tensor to decompose, since they are all of the same shape, allowing for a straightforward application of tensor decomposition techniques.
@@ 4.2 Results
4.2.1 One Layer at a Time
Table 1 compares the performance of the baseline, LASER, and TRAWL methods for both RoBERTa and GPT-J models in three data sets: BigBench WikiQA, BiosProfession, and HotpotQA.
First, we observe that TRAWL, our tensor decomposition approach across multiple weight matrices, consistently outperforms the baseline compression method, LASER. For the BigBench WikiQA dataset, TRAWLCP delivers the best accuracy, with 43.46% for RoBERTa and 68.1% for GPT-J, outperforming LASER and the baseline. It also achieves the lowest loss for GPT-J, while LASER has the lowest loss for RoBERTa. In the BiosProfession dataset, TRAWLCP continues to lead with 73.07% accuracy for RoBERTa and 82.35% for GPT-J, while improving the loss of GPT-J. For HotpotQA, the improvements are smaller, with TRAWLCP slightly ahead of LASER for GPT-J, and similar results between methods for RoBERTa.
Summary of results for one-layer-at-a-time: In this setting, TRAWL consistently outperforms the baseline and LASER methods, particularly with CP decomposition in the FC layers. The most significant improvements are seen with the GPT-J model, especially on the BigBench WikiQA and BiosProfession datasets. However, the observed benefits are dependent on the data set, and HotpotQA shows only marginal improvements.
4.2.2 Across the Entire Model
When we applied CP decomposition across all FC layers of the GPT-J model, we observed a significant drop in performance. As shown in Figure 4, the accuracy in the WikiQA dataset decreased when all layers were decomposed, a trend consistent between different ranks and components such as QKVO matrices. The combined decomposition likely introduced noise due to the varying characteristics of each layer, making it harder to capture the patterns needed to maintain performance. As a result, we switched to a layer-by-layer approach, which proved to be more effective.
Summary of results for across-model approach: The employed tensor formulation method did not yield promising outcomes across all datasets in the QKVO and FC cases. A possible explanation for this is that the latent patterns across all layers are widely heterogeneous and are not effectively captured within the rank budget that we set for our experiments. This may be potentially mitigated by using more complex tensor models that may be able to capture the heterogeneity more effectively. We defer this exploration to future work.
4.2.3 Segmented Layers
When organizing the chosen matrices into predefined sections and dividing them roughly into thirds for the initial, middle, and final parts of the model, we observed several different things.
For RoBERTa, we have our results for this tensor formulation shown in Table 2. The first thing we notice is that the highest gain in accuracy occurs when decomposing the last third of the FC weights on both WikiQA and BiosProfession datasets. Besides, the early and last segments of the QKVO matrices exhibit the most declines in performance. These results underscore the importance of focusing on the final segments of the FC layers for optimal performance improvements. They also suggest that QKVO matrices are less impactful in this tensor decomposition framework.
Summary of results for segmented layers: This approach only produces improved accuracies when decomposing the last FC matrices. For all the other combinations including early QKVO, middle QKVO, last QKVO, early FC, and middle FC, we get a drop in accuracy. Our results show that there is a smaller drop in accuracy when decomposing the middle third of the layers for QKVO weight matrices, indicating these layers are more compressible than the rest. In this type of tensor formulation, the highest gains were seen by stacking the FC matrices of the final few layers, although the gain was not as high as decomposing them one at a time.
| WikiQA | BiosProfession | ||
| QKVO | E | 0.17 | 30.49 |
| M | 27.11 | 51.06 | |
| L | 0.87 | 20.00 | |
| FC | E | 0.20 | 55.16 |
| M | 4.15 | 52.19 | |
| L | 31.53 | 63.41 | |
| Baseline Acc | 27.98 | 64.57 | |
@@ 4.3 Ablations
To evaluate the effect of stacking and decomposing different layers, we compared the performance of CP-decomposed fully connected (FC) weight matrices across various layers of RoBERTa on multiple datasets. The most notable improvements occurred in the FC weight matrices of the final layers, where tensor decomposition led to significant accuracy gains. In contrast, decomposing the QKVO matrices yielded minimal benefits and, in some cases, even reduced performance. As shown in Figure 3, the later-stage FC layers exhibited the largest accuracy improvements, while earlier layers, particularly the QKVO matrices, showed little to no benefit, and in some instances, decreased accuracy. We also observe similar trends on GPT-J, where the results are presented in Figure 5.
@@
Last but not least, the one-layer-at-a-time approach usually resulted in the highest performance gains near the final few layers when decomposing FC weights. However, no significant improvements were observed for QKVO matrices. These results align with the findings of [15], highlighting that FC layers are key targets for effective compression. In this work we demonstrate that there exists meaningful low-rank structure when treating the FC weights jointly as a tensor. In the future, we aim to explore whether more complex models or alternative methods can yield improvements for QKVO matrices.
@@ 5 Conclusion
In this study, we introduce TRAWL, a novel method for denoising LLMs using tensor decomposition. TRAWL extracts specific weight matrices, stacks them into a tensor, and applies decomposition to replace the original matrices with low-rank approximations. Our results show that layer-by-layer decomposition and selective layer stacking significantly improve performance, with the most notable gains seen in the fully connected layers of the terminal layers. These findings highlight the effectiveness of adaptive decomposition techniques in enhancing LLM efficiency, sometimes even outperforming original models by denoising.
Our research presents several promising avenues for future exploration. One area that we would like to explore is the investigation of novel decomposition techniques. Alternatives like Block-Term Decomposition and randomized decomposition methods might have the potential to enhance performance even further. Additionally, enhancing the rank selection procedure is crucial, as the optimal rank plays a pivotal role in achieving optimal performance. Furthermore, we propose to extend the applicability of TRAWL by implementing it in a wider range of models and tasks. This includes, but is not limited to, domains within computer vision, natural language processing, and reinforcement learning.
@@ 6 Limitations
Acknowledgments
Research was supported by the National Science Foundation under CAREER grant no. IIS 2046086, grant no. CNS 2106982, and CREST Center for Multidisciplinary Research Excellence in CyberPhysical Infrastructure Systems (MECIS) grant no. 2112650, by the Agriculture and Food Research Initiative Competitive Grant no. 2020-69012-31914 from the USDA National Institute of Food and Agriculture, and by the University Transportation Center for Railway Safety (UTCRS) at UTRGV through the USDOT UTC Program under Grant No. 69A3552348340.